OpenStudy (anonymous):

Simplify 3 over quantity of 5 minus 6 i quantity of 15 plus 18 i over 61 quantity of negative 15 plus 18 i over 61 3 i over quantity of 6 plus 5 i quantity of 15 plus 18 i over negative 11

mathslover (mathslover):

"3 over quantity of 5 minus 6i" can be written as : \(\cfrac{3}{5-6i}\) , right?

OpenStudy (anonymous):

correct.

mathslover (mathslover):

Okay, so, in order to simplify this, we will multiply numerator and denominator by the conjugate of 5-6i Can you tell me what is the conjugate of 5 - 6 i ?

OpenStudy (anonymous):

(2 + i)(3 + 4i)?

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

@kiaracakes120513 what about the first question? it is not yet solved..

OpenStudy (anonymous):

umm i have no clue..

mathslover (mathslover):

Do you know about conjugate of "a + bi" ?

OpenStudy (anonymous):

nope..

mathslover (mathslover):

See, the conjugate of a+bi is a-bi

mathslover (mathslover):

so, conjugate of 5 -6i is 5 + 6i

OpenStudy (anonymous):

ohhh okay so it basically you just switch the sign

mathslover (mathslover):

Right!

mathslover (mathslover):

So, we will multiply 5 + 6i to numerator and denominator ... \(\cfrac{3}{5-6i} \times \color{Blue}{\cfrac{5 + 6i}{5 + 6i} }\) \(\cfrac{3(5+6i)}{(5-6i)(5+6i)}\)

mathslover (mathslover):

Now we will use (a-b)(a+b) = a^2 - b^2 identity here for the denominator.

mathslover (mathslover):

Can you simplify it further?

OpenStudy (anonymous):

no..?

mathslover (mathslover):

Okay, we will first simplify the denominator... AS : \((a+b)(a-b) = a^2 - b^2 \) So , (5+6i)(5-6i) = ?

OpenStudy (anonymous):

61

mathslover (mathslover):

Yep... good! So, denominator becomes 61 and numerator now..

mathslover (mathslover):

3(5+6i) = 3(5) + 3(6i) = ?

OpenStudy (anonymous):

it simplifies out to 0

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

No... 3(5) + 3(6i) = 15 + 18 i

OpenStudy (anonymous):

Ohh

mathslover (mathslover):

So, you get (15+18i)/61

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so thats the answer?

mathslover (mathslover):

You need to convert this in statement now.

OpenStudy (anonymous):

what does that mean? lol

mathslover (mathslover):

Just like the question was written and we converted it into equations! So, you do vice versa..

OpenStudy (anonymous):

ohh okayy